OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (-1, 6) and has a slope of -3. y - 6 = -3x - 3 y - 6 = -3(x + 1) y = -3x + 3 3x + y = 3

OpenStudy (zpupster):

point-slope form y-y1=m(x-x1) plug in m=slope y-y1=-3(x-x1 ) ordered pair (x,y) or (-1,6) plug in y-6=-3( x-(-1) ) you see now?

OpenStudy (anonymous):

Thank you so much that really helped!! can you help me with a few more?

OpenStudy (zpupster):

just post all questions indivually it is better that way

OpenStudy (anonymous):

okay I have a few like 3 or 4

OpenStudy (zpupster):

sure i will look for them

OpenStudy (anonymous):

thanks first one is Given the equation y - 3 = one half(x + 6) in point-slope form, identify the equation of the same line in standard form. x - 2y = -12 y = one halfx y = one halfx + 6 y = one halfx + 9

OpenStudy (anonymous):

I am just trying to see if I got them right before I submit this. on this one I said D

OpenStudy (zpupster):

hmmm you first have to mult 1/2(x+6) that is 1/2x+3

OpenStudy (zpupster):

you see

OpenStudy (anonymous):

okay so this one is A?

OpenStudy (zpupster):

no then add 3 to both sides y= 1/2x+3+3

OpenStudy (anonymous):

yeah I am sorry for that I miss read what you typed. I said D because 6+3=9 and that was the only one with 9

OpenStudy (zpupster):

y=1/2x+6 so C

OpenStudy (anonymous):

OOOOHHHHHH.....Thanks I'll post the next one

OpenStudy (anonymous):

A computer manufacturer built a new facility for assembling computers. There were construction and new equipment costs. The company paid for these costs and made combined profits of $40 million after 4 years while profits increased $30 million per year. Select the correct graph of this function. First file is A second is B third is C(my answer) and fourth is D

OpenStudy (zpupster):

look at the graphs they started out negative and at uear 4 = 40 mil which graph/??

OpenStudy (anonymous):

okay I look and now I'm think a or d. because both have the top of the line at 4

OpenStudy (zpupster):

no d at year 4 is over 150 million we are looking for year 4 = 40 mil

OpenStudy (anonymous):

okay so A? Sorry Im not good at the ones with graphs as the answers

OpenStudy (zpupster):

no we are looking for year 4 = 40 mil that started at a profit year o is already has millions look at year 4

OpenStudy (anonymous):

okay I looked again and im gonna say b. because u said a,c,and d aren't it

OpenStudy (anonymous):

B or C? I look at what u said, u didnt say c wasnt it

OpenStudy (zpupster):

b

OpenStudy (anonymous):

okay thanks I have one more like this with graphs they are named the same as these but are different. THIS IS THE LAST QUESTION I HAVE TO CHECK IF ITS RIGHT OR NOT.

OpenStudy (anonymous):

A computer manufacturer built a new facility for assembling computers. There were construction and new equipment costs. The company paid for these costs and made combined profits of $80 million after 4 years while profits increased $60 million per year. Select the correct graph of this function. I said B. 1-A 2-B(my answer) 3-C 4-D